Monthly coffeecoffee cost​ ($) Number of students ​10.00-19.99 10 ​20.00-29.99 18 ​30.00-39.99 27 ​40.00-49.99 12 ​50.00-59.99...

Monthly

coffeecoffee

cost​ ($)

Number of students

​10.00-19.99

10

​20.00-29.99

18

​30.00-39.99

27

​40.00-49.99

12

​50.00-59.99

7

A sample of college students was asked how much they spent monthly on

coffee.

Approximate the standard deviation for the cost.

The sample standard deviation for the cost is.....

​(Round to the nearest cent as​ needed.)
Answer #1

Observation Table:

Cost Mid point Xi Freq . Fi xi * fi fi * Xi2
10 - 19.99 14.995 10 149.95 2248.5
20 - 29.99 24.995 18 449.91 11245.5
30 - 39.99 34.995 27 944.865 33065.55
40 - 49.99 44.995 12 539.94 24294.6
50 - 59.99 54.995 7 384.965 21171.15
74 2469.63 92025.3

Mean = = 2469.63 / 74 = 33.37

= 92025.3 / 74 - (33.37)2

= 130.03

Standard deviation =   = 11.40